OpenStudy (anonymous):

Laplace Transform of t ^{2}sin(2t)?

OpenStudy (jamesj):

Apply the same trick we just used twice.

OpenStudy (anonymous):

I have this table where I know the laplace transform of f(t)=t\[ t ^{n}*f(t)\] is \[ (-1)^n *F^{n}(s)\] where n=1,2,3 I'm just confused because the solution said that I can use that with n=2 OR n=1 to solve for the laplace transform above . n=2 makes sense since the function has a t^2 but I don't see why you can use n=1??

OpenStudy (jamesj):

To be clear, that's not the n-th power of F, that's the n-th derivative of F.

OpenStudy (anonymous):

Yeah I realized that earlier but why can you use n=1 because isn't the function t^2?

OpenStudy (jamesj):

because maybe you know the Laplace transform of t.sin(at)

OpenStudy (anonymous):

Yeah I do but where does the other t go?

OpenStudy (anonymous):

sorry I'm new to this stuff that's why all the questions

OpenStudy (jamesj):

You would use the case n = 1, if you knew F where F(s) = L[t.sin 2t]. Then the Laplace transform you want L[t^2.sin2t] = - F'(s) = (-1)^1 F^(1) (s)

OpenStudy (jamesj):

But I'm guessing you don't; I don't off the top of my head. But it doesn't matter of course because I do know G(s) = L[sin 2t] and then it's easy to calculate L[t^2.sin 2t] = G''(s)

OpenStudy (jamesj):

Or put another way, G'(s) = -L[t.sin 2t] = -F(s), hence G''(s) = -F'(s)

OpenStudy (anonymous):

Ohhh I see that make sense it just makes the calculation easier!! Yeah I do know that because it's in the table he provides

OpenStudy (anonymous):

thank you!!